OpenStudy (anonymous):

What is the greatest common factor of 24a^2b and 18abc

OpenStudy (anonymous):

What is the GCF of 24 and 18

OpenStudy (anonymous):

\[ 18 =2 \times 3^2\\ 24=2^3 \times 3 \]

OpenStudy (anonymous):

GCF(18,24)= 2(3) =6

OpenStudy (anonymous):

GCF(f 24a^2b and 18abc)=6 a b